OpenStudy (anonymous):

simplify the following expression and rewrite it in equivalent form with positive exponents. (-2x^-2)-^4

OpenStudy (anonymous):

um I think, (-2x^(1/2))^(1/4) not sure.

OpenStudy (anonymous):

nope totally wrong ignore that lmao

OpenStudy (anonymous):

= 1/ (-2x^-2)^4

OpenStudy (anonymous):

= 1/(-2 / x^2)^4 = 1 / 16/x^8 = x^8 / 16

OpenStudy (anonymous):

was the original expression = (-2x^-2)^-4 ?

OpenStudy (anonymous):

r u there?